Madge Mobiles Inc., manufactures plastic moldings for truck seats. Its costing system utilizes two cost categories, direct materials and conversion costs. Each product must pass through Department A and Department B. Direct materials are added at the beginning of production. Conversion costs are allocated evenly throughout production.

Data for Department A for February 2019 are:
   Work in process, beginning inventory, 40% converted   200 units
   Units started during February   800 units
   Work in process, ending inventory   100 units

Costs for Department A for February 2019 are:
   Work in process, beginning inventory:
   Direct materials   $150,000
   Conversion costs   $200,000
   Direct materials costs added during February   $600,000
   Conversion costs added during February   $425,000

a) What is the unit cost per equivalent unit of beginning inventory in Department A?

B) How many units were completed and transferred out of Department A during February?
Answer #1

a) Unit cost per equivalent unit of material = 150000/200 = 750

Unit Cost per equivalent unit of conversion = 200000/80 = 2500

Total Unit cost per equivalent unit of beginning inventory in Department A = 2500+750 = $3250

b) Unit completed and transferred out = 200+800-100 = 900 Units